OpenStudy (anonymous):

this is a tough one...... Simplify: 4 6/7 + 9 1/4 + 2 3/4 + 1/7.

OpenStudy (anonymous):

Why is it tough?

OpenStudy (anonymous):

because i suck at simplyfying

OpenStudy (anonymous):

Take the middle two ....

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

U know what 1/4 and 3/4 is...

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

Same game with the other 2 terms...

OpenStudy (anonymous):

see here: we have to just convert mixed fractions into improper fractions and then solve easily 4 6/7 = 34/7 9 1/4 = 37/4 2 3/4 = 11/4 1/7 =1/7 now see \[34/7 + 37/4 + 11/4 + 1/7\] \[34/7 +1/7 +11/4+37/4\] \[35/7 +48/4\] \[5+12\] \[17\]

OpenStudy (anonymous):

ookkk.....

OpenStudy (anonymous):

got it mathmaster

OpenStudy (anonymous):

just convert the mixed fractions into improper fractions then it would be easy

OpenStudy (anonymous):

\[34/7+37/4+11/4+1/7\]

OpenStudy (anonymous):

So 9+2 +1 and 4+1

OpenStudy (anonymous):

17...

OpenStudy (anonymous):

AIH KUSH YOURE SO SPEEDER THAN ME

OpenStudy (anonymous):